RC meeting and Run Plan, Friday, November 12th

From SBS wiki
Jump to: navigation, search

Daily_Meetings

RC Meeting Information

RC Daily Meetings are at 3:15pm, remote connection link: https://bluejeans.com/256205934

Accelerator Status

The RF separator work required to set up Hall A for 4th-pass beam was accomplished this morning in anticipation of the authorization for Hall A to resume operations.

See 3941723

Hall A Status and News

We have obtained the run certificate allowing us to resume the experiment. We will start at 4th pass and run here until the change to 5th pass on Nov. 24.

Updates from this meeting:

Hall Status:

  • Work in the Hall is finished. Abandoned the load test. Jack final testing of magnet power supply now. Re-test SBS magnet up to full current + correctors and BigBite as well.
  • Completed pre-beam checklist.
  • When Jack is done.

Subsystem Updates

  1. HCAL: ready for beam. Verify SBS magnet settings with LH2 elastic
  2. BBCal: One shower channel missing signal from last night's cosmic run. Loose cable connection. Request: need 1 hour cosmic run with magnet on.
  3. Timing hodoscope: no special needs, but may do a HV scan at some point.
  4. Grinch: nothing new, idea to take an LED run with BB on but not SBS to see the effect. ~1 hour.
  5. GEMS: No update on GEMs. Ramping and setting trip limits for beam.
  6. Beamline: No update.
  7. DAQ: Cosmics overnight ran stable.
  8. Replay analysis: Updated SBS-offline, standard replay appears to be working now on the CH machines.